OpenStudy (anonymous):

Solve the following equation analytically. 2 log7(x) = log7(2) + log7(2x + 16)

OpenStudy (anonymous):

\[\log_7(x^2)=\log_7(2(2x+16))\] \[x^2=4x+32\] \[x^2-4x-32=0\] \[(x-8)(x+4)=0\] \[x=8\] forget about \[x=-4\] because you cannot take the log of a negative number